# On Championing Foundation Models: From Explainability to Interpretability

### v1

@@ 5.3.2 Concluding Remarks
Some reason of hallucination is from limited expressive power: One theoretical study suggests that hallucination is unavoidable for FMs [52], which explains the occurrence of risks as illustrated in Figure 14 and 13. Though their definition of hallucination (existing gap between the model’s output and the unique ground truth function) is strict in some real-world scenarios where the correct answer isn’t always unique, e.g. Figure 14, it explains some types of hallucination, e.g. Figure 13. In these cases, it becomes crucial to enhance public awareness and develop strategies to minimize the impact of FMs’ hallucinations.
@@ 6 Conclusion and Future Research Direction
In this perspective review, we offer a comprehensive overview of interpretable techniques for FMs, encompassing generalization capability analysis, expressive power analysis, and dynamic behavior analysis. These interpretable methods illuminate the fundamental reasons behind various emerging properties in FMs, such as in-context learning, chain-of-thought reasoning, and potential risks like privacy breaches, biases, and hallucinations. As FMs advance, the need for interpretability becomes increasingly critical to guarantee transparency and trust-worthy. Furthermore, we delve into a captivating research frontier and promising directions for higher-level interpretation on these powerful FMs systems:
@@
  • Can we scale forever? -- All for AI: Recent theoretical studies suggest that increasing both the training dataset size and model complexity enhances the performance of FMs, specifically in terms of their generalization [176] and reasoning capabilities [93, 101]. The recent emergence of FMs, such as Llama-3.1 405B, has been hailed as another triumph of scaling laws. However, the immense data requirements and computational resources needed to train cutting-edge AI models have created significant barriers to entry for many researchers, startups, and smaller tech firms. Thus, the debate over the continued effectiveness of scaling is ongoing. While ‘scale is all you need’ seems mostly true for direct training, when it comes to transfer learning, the downstream performance critically depends on the tasks at hand as well as the choice of architecture and hyperparameters [177]. [178] show that some smaller models have achieved performance comparable to or even better than larger models. For instance, MiniCPM, a model with only 2 billion parameters, performs similarly to Llama, which has 13 billion parameters. Given the current situation, we should, on the one hand, conduct extensive theoretical and empirical research into the boundaries and limits of model scaling laws to proactively predict their performance growth and fully leverage the potential of FMs. On the other hand, we need to integrate high-performing small-scale models, develop effective collaborative learning frameworks, and reshape the AI landscape towards a future where “all is for AI”.
  • Self-consuming learning framework: Our interpretations collectively illustrate a phenomenon—high-quality data positively impacts model performance. However, as model complexity grows, the demand for data also escalates, yet a sobering reality is the increasing scarcity of high-quality data. In fact, the extensive repository of publicly accessible data on the internet has nearly been depleted [179]. To address this issue, many studies are dedicated to exploring self-consuming optimization frameworks, wherein generative models use simulated data to train themselves [180]. Recently, empirical assessments have indicated that the diminishing heavy-tail characteristics in the synthetic data distribution are a crucial factor affecting the efficacy of this optimization framework [181]. Hence, developing new interpretable method, theoretically constructing the generalization error of self-consuming optimization, and characterizing the dynamic changes in data distribution are vital strategies to further improve this optimization framework.

### v2

@@ 6.3.2 Concluding Remarks
Some reason of hallucination is from limited expressive power: One theoretical study suggests that hallucination is unavoidable for FMs [213], which explains the occurrence of risks as illustrated in Figure 16 and 15. Though their definition of hallucination (existing gap between the model’s output and the unique ground truth function) is strict in some real-world scenarios where the correct answer isn’t always unique, e.g. Figure 16, it explains some types of hallucination, e.g. Figure 15. In these cases, it becomes crucial to enhance public awareness and develop strategies to minimize the impact of FMs’ hallucinations.
@@ 7 Conclusion and Future Research Direction
In this survey, we offer a comprehensive overview of interpretable techniques for FMs, encompassing generalization capability analysis, expressive power analysis, and dynamic behavior analysis. These interpretable methods illuminate the fundamental reasons behind various emerging properties in FMs, such as in-context learning, chain-of-thought reasoning, and potential risks like privacy breaches, biases, and hallucinations. As FMs advance, the need for interpretability becomes increasingly critical to guarantee transparency and trust-worthy. Furthermore, we delve into a captivating research frontier and promising directions for higher-level interpretation on these powerful FMs systems:
@@
  • Can we scale forever? Recent theoretical studies suggest that increasing both the training dataset size and model complexity enhances the performance of FMs, specifically in terms of their generalization [220] and reasoning capabilities [86, 88]. The recent emergence of FMs, such as Llama-3.1 405B, has been hailed as another triumph of scaling laws. However, the immense data requirements and computational resources needed to train cutting-edge AI models have created significant barriers to entry for many researchers, startups, and smaller tech firms. Thus, the debate over the continued effectiveness of scaling is ongoing. While ‘scale is all you need’ seems mostly true for direct training, when it comes to transfer learning, the downstream performance critically depends on the tasks at hand as well as the choice of architecture and hyperparameters [221]. [222] show that some smaller models have achieved performance comparable to or even better than larger models. For instance, MiniCPM, a model with only 2 billion parameters, performs similarly to Llama, which has 13 billion parameters. Given the current situation, we should, on the one hand, conduct extensive theoretical and empirical research into the boundaries and limits of model scaling laws to proactively predict their performance growth and fully leverage the potential of FMs. On the other hand, we need to integrate high-performing small-scale models, develop effective collaborative learning frameworks, and reshape the AI landscape towards a future where “all is for AI”.
  • Does emergence truly exist? The emergence of FMs has brought about a new class of AI systems with remarkable capabilities. These models exhibit a range of emergent properties, e.g., zero-shot learning [7] and emergent capability [15], which go beyond their initial training objectives. Recent work [122] has proposed a quantifiable definition of emergence, describing it as a process where the entropy reduction of the entire sequence exceeds the entropy reduction of individual tokens. This offers a metric for evaluating emergent behaviors, although it remains limited in its generalizability across different tasks and models. However, the current understanding of emergent abilities is still in its infancy, particularly in terms of rigorous theoretical analysis of scaling laws. This reveals a gap in our knowledge of the relationship between model size and emergent capabilities. Traditional statistical learning theory tools, such as concentration inequalities and complexity measures (e.g., VC dimension, Rademacher complexity), are typically designed for simpler models and rely on strong assumptions about the data or learning task. These assumptions often break down in the context of FMs, rendering these tools inadequate for accurately characterizing and explaining emergent capabilities. To rigorously analyze the existence and triggering conditions of emergent capabilities, new theoretical frameworks are needed, such as phase transition analysis from the perspective of dynamic system.